OpenStudy (lyric.life):

Solve each equation for the variables a. 10(8x+4)-5(x+5)=165

OpenStudy (lyric.life):

so far I got 80x+4-5x+s5=165

OpenStudy (hafeda):

distribute and factor out the x and constant on the other side

OpenStudy (lyric.life):

25*

OpenStudy (lyric.life):

How did I do with that^?

OpenStudy (hafeda):

it should look like as follow : 80x +40 -5x -25 = 165 x(80-5) = 165-40+25

OpenStudy (hafeda):

distribute the 10 and 5 to both teams in the parenthesis

OpenStudy (lyric.life):

oh so you multiply 10 by 8 and 4 and 5 by x 5by 5 to get what u got then um subtract 80-5 to get 75=165-15?

OpenStudy (lyric.life):

for adding -40+25?

OpenStudy (hafeda):

75x = 165-15 yes you are write

OpenStudy (hafeda):

now divide both side by 75 to get x alone, but first finish the subtraction of 165-15

OpenStudy (hafeda):

opps sorry right lol

OpenStudy (lyric.life):

x=2?

OpenStudy (hafeda):

yes indeed now check it, just in case u got it wrong, to check just plug x = 2 in the equation

OpenStudy (lyric.life):

got it right ^-^ thank you for your time and help

OpenStudy (lyric.life):

can you help me with one more question? I'll open a new question

OpenStudy (lyric.life):

and close this one

OpenStudy (hafeda):

OK no problem